# Language Bottleneck Models: A Framework for Interpretable Knowledge Tracing and Beyond

### v1

@@ Abstract
Accurately assessing student knowledge is critical for effective education, yet traditional Knowledge Tracing (KT) methods rely on opaque latent embeddings, limiting interpretability. Even LLM-based approaches generate direct predictions or summaries that may hallucinate without any accuracy guarantees. We recast KT as an inverse problem: learning the minimum natural-language summary that makes past answers explainable and future answers predictable. Our Language Bottleneck Model (LBM) consists of an encoder LLM that writes an interpretable knowledge summary and a frozen decoder LLM that must reconstruct and predict student responses using only that summary text. By constraining all predictive information to pass through a short natural-language bottleneck, LBMs ensure that the summary contains accurate information while remaining human-interpretable. Experiments on synthetic arithmetic benchmarks and the large-scale Eedi dataset show that LBMs rival the accuracy of state-of-the-art KT and direct LLM methods while requiring orders-of-magnitude fewer student trajectories. We demonstrate that training the encoder with group-relative policy optimization, using downstream decoding accuracy as a reward signal, effectively improves summary quality.
@@ 1 Introduction
The Challenge of Estimating a Student’s Knowledge State.
@@
A fundamental objective in education is accurately assessing what a student knows, identifying misconceptions, and understanding how their knowledge evolves over time posner1982accommodation ; larkin2012misconceptions ; chen2020impact . Teachers intuitively achieve this through diagnostic reasoning: by observing students’ answers, they infer not merely correctness, but deeper patterns reflecting conceptual mastery or specific misunderstandings. For instance, a student consistently misapplying algebraic properties signals conceptual gaps requiring targeted intervention. Replicating this nuanced interpretability within AI would profoundly enhance personalized education at scale, augmenting human teaching by providing transparent, actionable insights into learners’ knowledge states (khosravi2022explainable, ).
Building algorithms that can learn the knowledge state of students as they interact with educational material is an active field of study known as Knowledge Tracing (KT) (corbett_knowledge_1994, ; shen2024survey, ) (see Appendix A for a detailed introduction to KT). However, accurately modeling a student’s knowledge state in a way that is both predictive and interpretable remains a significant open challenge. Traditional KT methods (piech_deep_2015, ; zhang2017dynamic, ; pandey2019self, ; ghosh2020context, ) rely primarily on statistical, embedding-based representations. Such models lack transparency, representing knowledge states as high-dimensional latent vectors that provide limited interpretability or actionable insights for teachers and learners (Figure 1B). A truly impactful AI assistant in education must go beyond opaque predictions, explicitly articulating what the student understands and what they do not, in an accurate and interpretable manner.
Limitations of Existing LLM-based approaches for KT.
@@
Recent approaches leveraging large language models (LLMs) have shown that LLMs can produce sensible predictions of students’ future behavior when provided with relevant information (li2024explainable, ; kim2024beyond, ), help mitigate the KT’s cold-start problem (lee2024language, ) and be finetuned to improve accuracy on KT tasks (wang2025llm, ). Nonetheless, these LLM-based methods still lack rigorous interpretability, as they either treat the model as a black-box or rely on free-form explanations that are susceptible to hallucination (bender2021dangers, ). Without ensuring faithful and grounded summaries, LLM-generated explanations fall short as reliable educational tools, emphasizing the need for a structured and principled foundation for generating interpretable, reliable, and adaptive student knowledge representations.
Figure 1: Language Bottleneck Models for Knowledge Tracing. The observed questions and answer $\mathcal{D}=\{\mathcal{X},\mathcal{Y}\}$ are caused by a certain knowledge state $\mathcal{S}$ held by the student when answering questions. The language bottleneck aims to approximate this knowledge state in natural language ($\tilde{\mathcal{S}}$). Then, the predicted knowledge state is used to predict both the observed historical data $\mathcal{X}$ and future questions $\mathcal{Y}$.
Knowledge Tracing as an Inverse Problem: Towards Concise and Interpretable Summaries.
Existing Knowledge Tracing methods typically approach student modeling as a forward prediction task, using latent statistical embeddings to predict future responses. However, this approach provides limited insight into the underlying student knowledge. We reframe Knowledge Tracing as an inverse problem: viewing student responses as outcomes of an underlying knowledge state $\mathcal{S}$, manifested through the student’s cognitive process (Figure 1A). The goal is to recover an interpretable representation $\tilde{\mathcal{S}}$ of the student knowledge state. Crucially, answers to downstream questions are expected to arise from the same latent knowledge state—an assumption we discuss further in Section 2.2. A useful representation must therefore reconstruct observed answers, enable prediction of future responses, and remain interpretable and steerable for teachers.
Introducing Language Bottleneck Models (LBMs).
Motivated by the inverse-problem framing described above, we introduce Language Bottleneck Models (LBMs). As illustrated in Figure 1C, LBMs uses two Large Language Models (LLMs) as student knowledge state extractor (encoder) and interpreter (decoder), generating succinct textual summaries meant to capture the student knowledge state. Specifically, LBMs map sequences of student interactions $\mathcal{X}$ to a concise, natural-language summary $\tilde{\mathcal{S}}$. This summary then serves as the sole basis for reconstructing the observed responses $\mathcal{X}$ and accurately predicting future responses $\mathcal{Y}$. By design, this ensures that the generated explanations encompass precisely the information necessary to interpret past behavior and reliably forecast future student performance.
Contributions and Paper Structure.
We make three key contributions:
  • We reframe knowledge tracing as an inverse problem and formalize a minimal-description objective that yields accurate and interpretable summaries of student knowledge states.
  • We introduce Language Bottleneck Models (LBMs)—an encoder-decoder architecture that forces all predictive information through a concise natural-language summary.
  • We empirically validate LBMs on synthetic and real-world datasets against strong KT and direct-prompting baselines, and show how reinforcement finetuning can train the encoder using downstream prediction accuracy.
The remainder of this paper is organized as follows: Section 2 introduces the knowledge tracing problem and reframes it as an inverse problem. Section 3 presents our Language Bottleneck Model (LBM) framework, along with results motivating this approach, and training methods. Section 4 discusses related literature on knowledge tracing and interpretability. Section 5 presents experimental results validating the effectiveness of LBMs. Finally, Section 6 discusses the applicability of LBMs beyond knowledge tracing and concludes with directions for future research.
2 Knowledge Tracing as an Inverse Problem
In this section, we formally introduce the knowledge tracing problem. We then reframe it as an inverse problem—inferring interpretable knowledge states from observed responses, outlining key desiderata for such representations and the conditions under which this formulation is appropriate.
2.1 Knowledge Tracing: Formal Problem Definition
Knowledge Tracing (KT) is fundamentally a sequence modeling problem where the objective is to infer a student’s latent knowledge state based on observed interactions. Formally, let $\mathcal{X}=\{(q_{1},a_{1}),(q_{2},a_{2}),\ldots,(q_{t},a_{t})\}$ denote a sequence of question-answer pairs, where $q_{i}\in\mathcal{Q}$ represents a question from the question space $\mathcal{Q}$, and $a_{i}\in\mathcal{A}$ represents the student’s answer from the answer space $\mathcal{A}$. The goal of KT is to learn a function that predicts the probability of correctly answering a future question $q_{t+1}$ given the historical interaction sequence: $P(a_{t+1}|q_{t+1},\mathcal{X})$.
While this formulation captures the predictive aspect of KT, it fails to address a crucial dimension: the interpretability. An ideal KT system should not only predict accurately but also provide educators with an explicit, human-understandable representation of the student’s knowledge state—such as mastered concepts, knowledge gaps and misconceptions. The importance of interpretability for KT is further discussed in the Appendix.
2.2 Knowledge Tracing as an Inverse Problem: From Black–Box Predictions to Interpretable Summaries
The conventional KT formulation (§2.1) focuses on predictive accuracy but offers limited insight into the underlying knowledge that drives student behavior. We instead frame KT as an inverse problem: in many practical scenarios—such as a quiz completed in a single sitting—a student’s responses can largely be explained by a fixed underlying knowledge state, expressed through the student’s cognitive process of answering questions. Most observed behavior reflects this latent state, with occasional deviations due to noise—such as slips (careless mistakes) or guesses (lucky correct answers), as formalized in Bayesian KT corbett_knowledge_1994 . The goal, then, is to recover an interpretable representation $\tilde{S}$ of the knowledge state that accounts for the observed interactions and supports accurate future predictions.
Framing KT in this way highlights three desiderata. (1) Self-consistency: the inferred summary $\tilde{S}$ should reconstruct the past by making the observed answers in $\mathbf{X}$ likely under a decoder. (2) Predictive sufficiency: $\tilde{S}$ must retain all information necessary to accurately predict future responses to unseen questions. (3) Human interpretability and steerability: the summary should be readable and editable by domain experts, enabling the diagnosis of student misconceptions and manual refinement or augmentation of $\tilde{S}$ to guide future predictions.
When is the inverse formulation appropriate? Although real learning is dynamic, many high–stakes diagnostic situations (end–of–unit quizzes, placement tests, or short tutoring sessions) span minutes rather than weeks. In such windows, it is reasonable to assume a quasi–static knowledge state. In this work, we explicitly focus on scenarios characterized by such quasi–static knowledge states. For longer horizons, the knowledge state might be better modeled as piecewise-static or smoothly evolving $S(t)$. We leave the extension of our framework to dynamically evolving knowledge states for future work.
@@ 2.3 Natural Language as the Interface
While a student’s knowledge state is inherently latent, natural language offers a flexible and interpretable medium for expressing it. Unlike fixed concept vocabularies, natural language accommodates open-ended reasoning patterns, misconceptions, and levels of mastery. In this work, we propose to use concise natural language summaries as proxies for student knowledge, designed to capture the latent cognitive state in a way that is both human-readable and predictive. The next section introduces a model family that operationalizes this idea.
@@ 3 Language Bottleneck Models
Figure 2: Zero-shot knowledge state encoding compared to ground truth. LLM models with different capabilities are prompted to write a summary of the knowledge state (top left panel) of a student, given 100 observed questions and answers provided in text. The ground truth knowledge state (top right panel) describing the student behavior has three main components: constructs mastered, constructs not mastered, and misconceptions. All three models capture the construct mastery correctly, but are not able to capture any misconception, beside o1 which notices the negative numbers misconception (bottom row). Bottom right symbols correspond to components in Figure 1.
@@ 3.1 Formal Definition
To tackle the inverse problem formulation of KT introduced in Section 2, we propose Language Bottleneck Models (LBMs): a two-stage architecture for inferring interpretable and accurate knowledge states from a student’s past interactions.
Let $X_{s}\subseteq X$ denote a subset of observed student responses, and let $Y_{s}\subseteq Y$ denote a set of future questions whose answers the model aims to predict. An LBM is defined by:
  • An encoder $f_{\theta}:X_{s}\rightarrow S$, which maps a history of interactions to a textual summary $S$ of the student’s implicit knowledge state.
  • A decoder $g:S\rightarrow\hat{Y}_{s}$, which uses only the textual summary to predict the student’s answers to new questions.
The explicit use of natural language as a bottleneck encourages the model to produce summaries that balance brevity with completeness. If the language summary omits essential details, predictions of future responses will suffer; conversely, if the summary is overly detailed or redundant, its utility for educators decreases. Thus, this design ensures the resulting textual summary is simultaneously concise, informative, and directly interpretable by humans.
The following sections motivate this approach and outline how LBMs can be trained effectively.
@@ 3.2 Motivating Observations
Before detailing our training procedure, we highlight two motivating observations that justify the LBM framework and its constraints.
Observation 1: Summarizing a knowledge state is challenging. Figure 2 compares summaries produced by different LLMs when prompted to describe a student’s knowledge state from 100 question-answer pairs from a synthetic dataset (see Section 5 for details). While all models capture high-level skill mastery, only one correctly identifies a misconception (errors with negative numbers) out of the four existing ones, illustrating that standard summarization approaches often miss crucial latent patterns.
Figure 3: Accuracy on synthetic dataset given ground truth knowledge state summaries.
Observation 2: Given a good summary, capable LLMs can decode with high fidelity. Figure 3 evaluates the performance of different decoder models when given access to the ground truth summary of the student’s latent knowledge state (an example of a ground-truth summary is shown in Figure 2). Stronger models like GPT-4o achieve nearly perfect accuracy ($98\%$), indicating that the bottleneck representation is indeed sufficient to drive effective downstream prediction—provided it encodes the right information.
Together, these observations motivate the inverse formulation: while generating summaries from raw interactions is hard, once obtained, these summaries are powerful inputs for reasoning and prediction. This validates our approach of learning encoders that optimize predictive utility via interpretable, compressed natural language representations.
@@ 3.3 Training Objective and Optimization
We train LBMs to produce informative and interpretable summaries of student knowledge states. As demonstrated above, strong decoders such as GPT-4o can achieve near-perfect accuracy when provided with accurate knowledge summaries. We therefore optimize the encoder $f_{\theta}$ to produce such summaries, while freezing the decoder $g$ to a strong LLM during training.
Summary generation and evaluation.
Given a student’s interaction history $X$, we sample a subset $X_{\text{enc}}\subseteq X$ of $N$ interactions to generate a summary $S=f_{\theta}(X_{\text{enc}})$. We evaluate this summary on two disjoint subsets: (i) $X_{s}\subseteq X_{\text{enc}}$: past questions for assessing reconstruction ability; and (ii) $Y_{s}\subseteq Y$: future questions for evaluating predictive accuracy. The summary $S$ is passed to the frozen decoder $g$, which outputs predictions $\hat{X}_{s}=g(S;X_{s})$ and $\hat{Y}_{s}=g(S;Y_{s})$.
Reward function.
The reward for a summary $S$ is defined using a function $\Phi$ that balances multiple criteria:
$\mathcal{R}(S;g)=\Phi\left(\texttt{Acc}(\hat{X}_{s},X_{s}),\;\texttt{Acc}(\hat{Y}_{s},Y_{s}),\;|S|,\;\Omega(S)\right),$ (1)
where $\texttt{Acc}(\cdot)$ measures accuracy, $|S|$ is the summary length, and $\Omega(S)$ enforces structural constraints (e.g., inclusion of a Misconceptions section). The function $\Phi$ balances these components using hyperparameters or indicator functions to enforce constraints as needed.
Optimization via GRPO.

### v2

# Language Bottleneck Models for Qualitative Knowledge State Modeling

@@ Abstract
Accurately assessing student knowledge is central to education. Cognitive Diagnosis (CD) models estimate student proficiency at a fixed point in time, while Knowledge Tracing (KT) methods model evolving knowledge states to predict future performance. However, existing approaches either provide quantitative concept mastery estimates with limited expressivity (CD, probabilistic KT) or prioritize predictive accuracy at the cost of interpretability (deep learning KT). We propose Language Bottleneck Models (LBMs), where an encoder LLM produces textual knowledge state summaries, which a decoder LLM uses to predict future performance. This produces interpretable summaries that can express nuanced insights—such as misconceptions—that CD and KT models cannot capture. Extensive validation across synthetic and real-world datasets shows LBMs reveal qualitative insights beyond what CD and KT models can capture, while achieving competitive accuracy with improved sample efficiency. We demonstrate that the encoder and decoder can be fine-tuned with reinforcement learning and supervised fine-tuning respectively to improve both summary quality and predictive performance.
@@ 1 Introduction
Knowledge state modeling
@@
A fundamental objective in education is accurately assessing what a student knows, identifying misconceptions, and understanding how their knowledge evolves over time (Posner et al., 1982; Larkin, 2012; Chen et al., 2020). Teachers intuitively achieve this through diagnostic reasoning: by observing students’ answers, they infer not merely correctness, but deeper patterns reflecting conceptual mastery or specific misunderstandings.
Limitations of Cognitive Diagnosis and Knowledge Tracing
Cognitive Diagnosis (CD) (Templin et al., 2010; Wang et al., 2024) models produce diagnostic reports but are constrained to proficiency estimates over a fixed set of knowledge concepts. In parallel, Knowledge Tracing (KT) (Shen et al., 2024) models excel at predicting future performance based on observed past responses, yet remain limited to either estimating mastery over knowledge concepts (Corbett and Anderson, 1994; Zhou et al., 2024) or operating as black-box models requiring post-hoc interpretability (e.g. DKT (Piech et al., 2015; Ghosh et al., 2020)).
Limitations of existing LLM-based approaches
@@
Recent approaches leveraging large language models (LLMs) have shown that LLMs can produce sensible predictions of students’ future behavior when provided with relevant information (Li et al., 2024a; Kim et al., 2024), help mitigate the KT’s cold-start problem (Lee et al., 2024) and be finetuned to improve accuracy on KT tasks (Wang et al., 2025). Nonetheless, these LLM-based methods still lack rigorous interpretability, as they either treat the model as a black-box or rely on free-form explanations that are susceptible to hallucination (Bender et al., 2021). As a result, such methods fail to provide grounded, reliable representations of knowledge that can be trusted in educational practice. This gap motivates a principled approach where interpretability is an intrinsic design constraint.
Figure 1: Language Bottleneck Models for Knowledge Modeling. (A) Past and future behavior $\mathcal{X}$ and $\mathcal{Y}$ are caused by a certain knowledge state $\mathcal{S}$ held by the student when answering questions. (B) CD and KT models represent the knowledge state via quantitative proficiency vectors or opaque latent embeddings. (C) LBMs approximate the knowledge state using natural language summaries. The example knowledge states shown are taken from the case-study presented in Section 4.1.
Knowledge state modeling as an inverse problem
Assessing student knowledge can be framed as an inverse problem (Figure 1A): observable answers are generated by an underlying knowledge state $\mathcal{S}$, and the goal is to infer a faithful approximation $\tilde{\mathcal{S}}$. This formulation is commonly done in Cognitive Diagnosis, which constrains $\tilde{\mathcal{S}}$ to fixed quantitative proficiency estimates over knowledge concepts (Figure 1B), providing interpretability but lacking flexibility to capture nuanced knowledge states. In this work, we adopt an inverse-problem framing in which $\tilde{\mathcal{S}}$ takes the form of concise, free-form natural language. This shifts the knowledge state from proficiency estimates over knowledge concepts to a flexible description that can capture nuanced insights such as misconceptions.
Language Bottleneck Models
We introduce Language Bottleneck Models (LBMs), a modeling framework that uses natural language as an interpretable intermediate representation. Constraining all predictive information to flow through a textual bottleneck ensures that the model’s predictions are grounded in human-readable descriptions. In our knowledge state modeling setting (Figure 1C), an encoder LLM summarizes a student’s interaction history $\mathcal{X}$ into natural language $\tilde{\mathcal{S}}$, while a decoder LLM reconstructs past responses and predicts future ones using only that summary. This constitutes a language-based approach to Cognitive Diagnosis that captures nuanced insights beyond quantitative proficiency scores while maintaining strong predictive performance.
Contributions.
  • We cast knowledge state modeling as an inverse problem over open-ended textual representations—building on ideas from Cognitive Diagnosis but replacing quantitative concept proficiencies with a flexible natural-language knowledge state.
  • To instantiate this, we introduce Language Bottleneck Models (LBMs), which encode observed student behavior into predictive text-based summaries of knowledge states.
  • We present a rigorous qualitative evaluation using LLM-as-a-judge metrics and a detailed case study, demonstrating that LBMs provide nuanced insights—such as misconceptions—beyond what CD and KT models can capture.
  • We evaluate LBMs against 14 KT and CD baselines on three datasets, showing competitive accuracy with superior sample efficiency, and demonstrate that fine-tuning the encoder and decoder enhances predictive performance.
2 Knowledge State Modeling as an Inverse Problem over Textual Representations
2.1 Preliminaries and Notation
We consider data consisting of student interactions with educational questions. At each time step $t$, a student is presented with a question $q_{t}\in\mathcal{Q}$ and optionally knowledge concept (KC) information $k_{t}\in\mathcal{K}$ and provides a response $r_{t}\in\mathcal{R}$, which is evaluated for correctness $c_{t}\in\{0,1\}$. We represent an interaction as $x_{t}=(q_{t},k_{t},r_{t},c_{t})$, and a student’s interaction history up to time $t$ as $H_{t}=(x_{1},\ldots,x_{t})$.
The standard predictive task, often referred to as Knowledge Tracing (KT) (Corbett and Anderson, 1994; Shen et al., 2024), is to estimate $p(c_{t+1}\mid q_{t+1},H_{t})$, the probability that the student will answer a new question $q_{t+1}$ correctly, conditioned on their interaction history. This task definition captures the forward-prediction aspect of knowledge state modeling, but does not yet address how the underlying knowledge that drives these responses should be represented.
2.2 Knowledge State Modeling as an Inverse Problem
An alternative view is to frame knowledge state modeling as an inverse problem: observed responses are generated by a latent knowledge state $\mathcal{S}$ through the student’s cognitive process, and the goal is to recover an approximation $\tilde{\mathcal{S}}$ of this state from the observed responses.
This perspective is central to Cognitive Diagnosis (CD) models (Reckase, 2006; De La Torre, 2009; Templin et al., 2010; Wang et al., 2022, 2024), which produce diagnostic reports of concept mastery from observed responses. However, CD typically restricts $\tilde{\mathcal{S}}$ to quantitative mastery or proficiency vectors based on predefined or inferred concepts, limiting expressivity. Meanwhile, deep learning-based KT methods prioritize predictive accuracy without explicitly recovering $\tilde{\mathcal{S}}$, representing knowledge states as high-dimensional embeddings that lack transparency (Piech et al., 2015; Zhang et al., 2017; Pandey and Karypis, 2019; Ghosh et al., 2020). Bayesian KT (Corbett and Anderson, 1994) as well as recent KT extensions introduce diagnostic reports similar in spirit to CD, or use interpretable latent states (Yeung, 2019; Minn et al., 2022; Chen et al., 2023; Park et al., 2024; Zhou et al., 2024). However these approaches remain bound to estimates of KC mastery or rely on post-hoc interpretability.
Key assumption: constant knowledge state A practical assumption underlying this formulation is that a knowledge state can be treated as constant within short diagnostic windows (e.g., unit tests, placement exams, or tutoring sessions). This aligns with CD models (see §2.1 in Wang et al. (2024)), and distinguishes them from KT approaches which model the evolution of $\mathcal{S}_{t}$ across longer periods where the underlying knowledge state is expected to change.
@@ 2.3 Natural Language as the Interface
Our formulation follows CD approaches in adopting the inverse problem framing of inferring a diagnostic report from observed responses, but instead of restricting $\tilde{\mathcal{S}}$ to quantitative mastery scores, we model it through concise textual summaries. Natural language provides an interpretable and expressive medium—capable of describing arbitrary reasoning patterns or misconceptions, a key focus in education research (Smith III et al., 1994; Wang et al., 2020; King et al., 2024)—while remaining human-understandable. In the next section, we introduce Language Bottleneck Models (LBMs), which operationalizes this idea by compressing student interaction histories into concise textual representations that preserve predictive information.
@@ 3.1 Formal Definition
We propose Language Bottleneck Models (LBMs) for Knowledge State Modeling via textual summaries: an LLM-based, two-stage architecture designed to infer a predictive text-based knowledge state from a student’s interaction history.
[…]
In later sections, we describe how both the encoder $f_{\theta}$ and decoder $g_{\phi}$ can be trained. However, we first show with the following motivating experiments that encoding is a harder task than decoding: when given high-quality knowledge state summaries, strong off-the-shelf LLMs can achieve near-perfect prediction accuracy.
@@ 3.2 Motivating Observations
We motivate the design of LBMs by two observations.
Observation 1: Given a good knowledge state summary, strong LLMs can decode with high fidelity. To test this in an idealized setting, we used a Synthetic dataset where each student’s knowledge state and their resulting answers to questions are programmatically generated. Figure 2 evaluates the performance of different decoder models when given direct access to this perfect, "ground-truth" summary of the student’s latent knowledge (example knowledge state and summaries are shown in Figure A1 in the Appendix, and full dataset details are in §4). Stronger models like GPT-4o achieve nearly perfect accuracy ($98\%$), indicating that the bottleneck representation is indeed sufficient to drive effective downstream prediction for closed-form questions—provided it captures the right information.
Figure 2: Accuracy ± SEM (N=200) on Synthetic dataset given ground truth knowledge state summaries.
Observation 2: Summarizing knowledge states from raw interactions is non-trivial. Standard LLM summarization approaches can capture high-level skill mastery but often fail to identify crucial latent patterns like student misconceptions (see an example comparison on our Synthetic dataset Figure A1 in the Appendix).
Together, these observations suggest that the key challenge lies in learning an encoder that produces faithful summaries, rather than in the decoding step itself.
@@ 3.3 Training Objective and Optimization
3.3.1 Encoder
We propose a reinforcement learning-based approach to train an encoder to produce more faithful and predictive summaries by using downstream decoder accuracy as reward.
Summary generation and prediction Given an interaction history $H_{t}=(x_{1},\ldots,x_{t})$, the encoder $f_{\theta}$ maps a subset of interactions $\mathcal{X}_{\text{enc}}\subseteq H_{t}$ to a textual summary $\tilde{\mathcal{S}}=f_{\theta}(\mathcal{X}_{\text{enc}})$. The decoder $g$ then conditions on $\tilde{\mathcal{S}}$ to predict responses for two sets $\mathcal{X}$ and $\mathcal{Y}$: interactions used for reconstruction and prediction, respectively. In practice, $\mathcal{X}$ and $\mathcal{Y}$ may be chosen flexibly to include held-out past responses, future responses, or both.
Reward function Given predicted interactions $\tilde{\mathcal{X}}=\{g(q,\tilde{\mathcal{S}}),q\in\mathcal{X}\}$ and $\tilde{\mathcal{Y}}=\{g(q,\tilde{\mathcal{S}}),q\in\mathcal{Y}\}$, the reward for a summary $\tilde{\mathcal{S}}$ is defined as
$),R(\tilde{\mathcal{S}};g)=\phi\!\Big(\texttt{Acc}\big(\tilde{\mathcal{X}},\mathcal{X}\big),\texttt{Acc}\big(\tilde{\mathcal{Y}},\mathcal{Y}\big),\;|\tilde{\mathcal{S}}|,\;\Omega(\tilde{\mathcal{S}}$$\Big)$, (1)
[…]